# Adversarial Training Should Be Cast as a Non-Zero-Sum Game

### v1

@@ 5 Experiments
In this section, we evaluate the performance of BETA and BETA-AT on CIFAR-10 Krizhevsky et al. (2009). Throughout, we consider a range of AT algorithms, including PGD Madry et al. (2018), FGSM Goodfellow et al. (2015), TRADES Zhang et al. (2019), MART Wang et al. (2020), as well as a range of adversarial attacks, including APGD and AutoAttack Croce and Hein (2020). We consider the standard perturbation budget of $\epsilon=8/255$, and all training and test-time attacks use a step size of $\alpha=2/255$. For both TRADES and MART, we set the trade-off parameter $\lambda=5$, which is consistent with the original implementations Wang et al. (2020); Zhang et al. (2019).
The bilevel formulation eliminates robust overfitting. Robust overfitting occurs when the robust test accuracy peaks immediately after the first learning rate decay, and then falls significantly in subsequent epochs as the model continues to train Rice et al. (2020). This is illustrated in Figure 0(a), in which we plot the learning curves (i.e., the clean and robust accuracies for the training and test sets) for a ResNet-18 He et al. (2016) model trained using 10-step PGD against a 20-step PGD adversary. Notice that after the first learning rate decay step at epoch 100, the robust test accuracy spikes, before dropping off in subsequent epochs. On the other hand, BETA-AT does not suffer from robust overfitting, as shown in Figure 0(b). We argue that this strength of our method is a direct result of our bilevel formulation, in which we train against a proper surrogate for the adversarial classification error.
@@
BETA matches the robustness estimate of AutoAttack. AutoAttack is a state-of-the-art adversarial attack which is widely used to estimate the robustness of trained models on leaderboards such as RobustBench Croce et al. (2020a); Croce and Hein (2020). In brief, AutoAttack comprises a collection of four disparate attacks: APGD-CE, APGD-T, FAB, and Square Attack. AutoAttack also involves several heuristics, including multiple restarts and variable stopping conditions. In Table 2, we compare the performance of the top-performing models on RobustBench against AutoAttack, APGD-T, and BETA with RMSprop. Both APGD-T and BETA used thirty steps, whereas we used the default implementation of AutoAttack, which runs for 100 iterations. We also recorded the gap between AutoAttack and BETA. Notice that the 30-step BETA—a heuristic-free algorithm derived from our bilevel formulation of AT—performs almost identically to AutoAttack, despite the fact that AutoAttack runs for significantly more iterations and uses five restarts, which endows AutoAttack with an unfair computational advantage. That is, excepting for a negligible number of samples, BETA matches the robustness estimate of AutoPGD-targeted and AutoAttack, despite using an off-the-shelf optimizer.
@@ 6 Related work
Robust overfitting. Several recent papers (see, e.g., (Rebuffi et al., 2021; Chen et al., 2021; Yu et al., 2022; Dong et al., 2022; Wang et al., 2020; Lee et al., 2020)) have attempted to explain and resolve robust overfitting (Rice et al., 2020). However, none of these works point to a fundamental limitation of adversarial training as the cause of robust overfitting. Rather, much of this past work has focused on proposing heuristics for algorithms specifically designed to reduce robust overfitting, rather than to improve adversarial training. In contrast, we posit that the lack of guarantees of the zero-sum surrogate-based AT paradigm Madry et al. (2018) is at fault, as this paradigm is not designed to maximize robustness with respect to classification error. And indeed, our empirical evaluations in the previous section confirm that our non-zero-sum formulation eliminates robust overfitting.
Estimating adversarial robustness. There is empirical evidence that attacks based on surrogates (e.g., PGD) overestimate the robustness of trained classifiers Croce and Hein (2020); Croce et al. (2020b); Gowal et al. (2019). Indeed, this evidence served as motivation for the formulation of more sophisticated attacks like AutoAttack Croce and Hein (2020), which empirically tend to provide more accurate estimates of robustness. In contrast, we provide solid, theoretical evidence that commonly used attacks overestimate robustness due to the misalignment between standard surrogate losses and the adversarial classification error. Moreover, we show that optimizing the BETA objective with a standard optimizer (e.g., RMSprop) achieves the same robustness as AutoAttack without employing ad hoc training procedures such as multiple restarts. convoluted stopping conditions, or adaptive learning rates.
@@
Bilevel formulations of AT. Prior to our work, (Zhang et al., 2022) proposed a different pseudo-bilevel [^3] In a strict sense, the formulation of Zhang et al. (2022) is not a bilevel problem. In general, the most concise way to write a bilevel optimization problem is $\min_{\theta}f(\theta,\delta^{\star}(\theta))$ subject to $\delta^{\star}(\theta)\in\operatorname*{arg\,max}g(\theta,\delta)$. In such problems the value $\delta^{\star}(\theta)$ only depends on $\theta$, as the objective function $g(\theta,\cdot)$ is then uniquely determined. This is not the case in (Zhang et al., 2022, eq. (7)), where an additional variable $z$ appears, corresponding to the random initialization of Fast-AT. Hence, in (Zhang et al., 2022) the function $g(\theta,\cdot)$ is not uniquely defined by $\theta$, but is a random function realized at each iteration of the algorithm. Thus, it is not a true bilevel optimization problem in the sense of the textbook definition (Bard, 2013). formulation for AT, wherein the main objective was to justify the Fast AT algorithm introduced in (Wong et al., 2020). More specifically, the formulation in (Zhang et al., 2022) is designed to produce solutions that coincide with the iterates of Fast AT by linearizing the attacker’s objective. In contrast, our bilevel formulation appears naturally following principled relaxations of the intractable classification error AT formulation. In this way, the formulation in Zhang et al. (2022) applies only in the context of Fast AT, whereas our formulation deals more generally with the task of adversarial training.
@@ 7 Conclusion
In this paper, we rigorously studied the standard zero-sum formulation of adversarial training. We argued that the surrogate-based relaxation commonly employed to improve the tractability of this problem voids guarantees on the ultimate robustness of trained classifiers, resulting in weak adversaries and ineffective AT algorithms. This shortcoming motivated the formulation of a novel, yet natural bilevel approach to adversrial training and evaluation. In our paradigm, the adversary and defender optimize separate objectives, which constitutes a non-zero-sum game that preserves guarantees on robustness. Based on this formulation, we developed a new adversarial attack algorithm—BETA, which stands for BEst Adversarial Attack—and a concomitant AT algorithm, which we call BETA-AT. In our experiments, we showed that BETA-AT eliminates robust overfitting, which we argued is a direct result of optimizing an objective which is aligned with the goal of finding true adversarial examples. We also showed that even when early stopping based model selection is used, BETA-AT performed comparably to AT. And finally, we showed that BETA provides almost identical estimates of robustness to AutoAttack, indicating that when the adversarial objective closely matches the true objective, one need not resort to heuristics like multiple restarts, variable stopping conditions, and adaptive learning rate schedules to accurately estimate robustness.
With regard to the bilevel formulation in this paper, future directions abound. One could imagine applying this framework to other changes in the data space, including the kinds of distribution shifts that are common in fields like domain adaptation and domain generalization. A convergence analysis of BETA and an analysis of the sample complexity of BETA-AT are two more directions that we leave for future work. The prospect of applying more sophisticated bilevel optimization algorithmic techniques to this problem is also a promising avenue for future research.
@@ Acknowledgements
FL is funded through a Ph.D. fellowship awarded by the Swiss Data Science Center, a joint venture between EPFL and ETH Zurich. AR is funded through a fellowship from Amazon AWS in conjunction with the ASSET center for secure and trustworthy AI at the University of Pennsylvania. This work was supported by the Swiss National Science Foundation under grant number 200021_205011 and by the NSF-Simons Foundation’s Mathematical and Scientific Foundations of Deep Learning (MoDL) program on Transferable, Hierarchical, Expressive, Optimal, Robust, Interpretable NETworks (THEORINET).
@@ Appendix A Proof of proposition 1
Suppose there exists $\hat{\eta}$ satisfying $\left|\left|\hat{\eta}\right|\right|\leq\epsilon$ such that for some $j\in[K]$, $j\neq Y$ we have $M_{\theta}(X+\hat{\eta},Y)_{j}>0$, i.e., assume
@@
for such $\hat{\eta}$ and such $j$ we have $f_{\theta}(X+\hat{\eta})_{j}>f_{\theta}(X+\hat{\eta})_{Y}$ and thus $\operatorname*{arg\,max}_{j\in[K]}f_{\theta}(X+\hat{\eta})_{j}\neq Y$. Hence, such $\hat{\eta}$ induces a misclassification error i.e.,
$\hat{\eta}\in\operatorname*{arg\,max}_{\eta:\|\eta\|_{2}\leq\epsilon}\left\{\operatorname*{arg\,max}_{j\in[K]}f_{\theta}(X+\eta)_{j}\neq Y\right\}$ (24)
In particular if
$(j^{\star},\eta^{\star})\in\operatorname*{arg\,max}_{j\in[K]-\{Y\},\>\eta:\|\eta\|\leq\epsilon}M_{\theta}(X+\eta,Y)_{j}\Rightarrow\eta^{\star}\in\operatorname*{arg\,max}_{\eta:\|\eta\|_{2}\leq\epsilon}\left\{\operatorname*{arg\,max}_{j\in[K]}f_{\theta}(X+\eta)_{j}\neq Y\right\}$ (25)
Otherwise, assume
@@
then for all $\eta:\>\left|\left|\eta\right|\right|<\epsilon$ and all $j\neq Y$ we have $f_{\theta}(X+\eta)_{j}<f_{\theta}(X+\eta)_{Y}$, so that $\operatorname*{arg\,max}_{j\in[K]}f_{\theta}(x+\eta)_{j}=Y$ i.e., there is no adversarial example in the ball. In this case for any $\eta$, in particular In particular if
@@ Appendix B Smooth reformulation of the lower level
First, note that the problem in eqs. 20, 21 and 22 is equivalent to
In this formulation, both upper- and lower-level problems are smooth (barring the possible use of nonsmooth components like ReLU). Most importantly (I) the smoothing is obtained through a lower bound of the original objective in eqs. 21 and 22, retaining guarantees that the adversary will increase the misclassification error and (II) all the adversarial perturbations obtained for each class now appear in the upper level (38), weighted by their corresponding negative margin. In this way, we make efficient use of all perturbations generated: if two perturbations from different classes achieve the same negative margin, they will affect the upper-level objective in fair proportion. This formulation gives rise to algorithm 3.

### v2

@@ 5 Experiments
In this section, we evaluate the performance of BETA and BETA-AT on CIFAR-10 (Krizhevsky et al., 2009). Throughout, we consider a range of AT algorithms, including PGD (Madry et al., 2018), FGSM (Goodfellow et al., 2015), TRADES (Zhang et al., 2019), MART (Wang et al., 2020), as well as a range of adversarial attacks, including APGD and AutoAttack (Croce & Hein, 2020). We consider the standard perturbation budget of $\epsilon=8/255$, and all training and test-time attacks use a step size of $\alpha=2/255$. For both TRADES and MART, we set the trade-off parameter $\lambda=5$, which is consistent with the original implementations (Wang et al., 2020; Zhang et al., 2019).
The bilevel formulation eliminates robust overfitting. Robust overfitting occurs when the robust test accuracy peaks immediately after the first learning rate decay, and then falls significantly in subsequent epochs as the model continues to train (Rice et al., 2020). This is illustrated in Figure 0(a), in which we plot the learning curves (i.e., the clean and robust accuracies for the training and test sets) for a ResNet-18 (He et al., 2016) trained using 10-step PGD against a 20-step PGD adversary. Notice that after the first learning rate decay at epoch 100, the robust test accuracy spikes, before dropping off in subsequent epochs. On the other hand, BETA-AT does not suffer from robust overfitting, as shown in Figure 0(b). We argue that this strength of our method is a direct result of our bilevel formulation, in which we train against a proper surrogate for the adversarial classification error.
@@
BETA matches the performance of AutoAttack. AutoAttack is a state-of-the-art attack which is widely used to estimate the robustness of trained models on leaderboards such as RobustBench (Croce et al., 2020a; Croce & Hein, 2020). In brief, AutoAttack comprises a collection of four disparate attacks: APGD-CE, APGD-T, FAB, and Square Attack. AutoAttack also involves several heuristics, including multiple restarts and variable stopping conditions. In Table 2, we compare the performance of the top-performing models on RobustBench against AutoAttack, APGD-T, and BETA with RMSprop. Both APGD-T and BETA used thirty steps, whereas we used the default implementation of AutoAttack, which runs for 100 iterations. We also recorded the gap between AutoAttack and BETA. Notice that the 30-step BETA—a heuristic-free algorithm derived from our bilevel formulation of AT—performs almost identically to AutoAttack, despite the fact that AutoAttack runs for significantly more iterations and uses five restarts, which endows AutoAttack with an unfair computational advantage. That is, excepting for a negligible number of samples, BETA matches the performance of AutoPGD-targeted and AutoAttack, despite using an off-the-shelf optimizer.
@@ 6 Related work
Robust overfitting. Several recent papers (see, e.g., (Rebuffi et al., 2021; Chen et al., 2021; Yu et al., 2022; Dong et al., 2022; Wang et al., 2020; Lee et al., 2020)) have attempted to explain and resolve robust overfitting (Rice et al., 2020). However, none of these works point to a fundamental limitation of AT as the cause of robust overfitting. Rather, much of this past work has focused on proposing heuristics for algorithms specifically designed to reduce robust overfitting, rather than to improve AT. In contrast, we posit that the lack of guarantees of the zero-sum surrogate-based AT paradigm Madry et al. (2018) is at fault, as this paradigm is not designed to maximize robustness with respect to classification error. And indeed, our empirical evaluations in the previous section confirm that our non-zero-sum formulation eliminates robust overfitting.
Estimating adversarial robustness. There is empirical evidence that attacks based on surrogates (e.g., PGD) overestimate the robustness of trained classifiers (Croce & Hein, 2020; Croce et al., 2020b). Indeed, this evidence served as motivation for the formulation of more sophisticated attacks like AutoAttack (Croce & Hein, 2020), which tend to provide more accurate estimates of robustness. In contrast, we provide solid, theoretical evidence that commonly used attacks overestimate robustness due to the misalignment between standard surrogate losses and the adversarial classification error. Moreover, we show that optimizing the BETA objective with a standard optimizer (e.g., RMSprop) achieves the same robustness as AutoAttack without employing ad hoc training procedures such as multiple restarts. convoluted stopping conditions, or adaptive learning rates.
@@
Bilevel formulations of AT. Prior to our work, (Zhang et al., 2022) proposed a different pseudo-bilevel [^3] In a strict sense, the formulation of (Zhang et al., 2022) is not a bilevel problem. In general, the most concise way to write a bilevel optimization problem is $\min_{\theta}f(\theta,\delta^{\star}(\theta))$ subject to $\delta^{\star}(\theta)\in\operatorname*{arg\,max}g(\theta,\delta)$. In such problems the value $\delta^{\star}(\theta)$ only depends on $\theta$, as the objective function $g(\theta,\cdot)$ is then uniquely determined. This is not the case in (Zhang et al., 2022, eq. (7)), where an additional variable $z$ appears, corresponding to the random initialization of Fast-AT. Hence, in (Zhang et al., 2022) the function $g(\theta,\cdot)$ is not uniquely defined by $\theta$, but is a random function realized at each iteration of the algorithm. formulation for AT, wherein the main objective was to justify the FastAT algorithm introduced in (Wong et al., 2020). Specifically, the formulation in (Zhang et al., 2022) is designed to produce solutions that coincide with the iterates of FastAT by linearizing the attacker’s objective. In contrast, our bilevel formulation appears naturally following principled relaxations of the intractable classification error AT formulation. In this way, the formulation in (Zhang et al., 2022) applies only in the context of Fast AT, whereas our formulation deals more generally with the task of AT.
In the same spirit as our work, (Mianjy & Arora, 2024) solve a problem equivalent to a bilevel problem wherein the adversary maximizes a “reflected” cross-entropy loss. While this paper focuses on binary classification, the authors show that this approach leads to improved adversarial robustness and admits convergence guarantees. Our approach, while related, is distinct in its reformulation of the adversarial training problem via the negative margin loss. Moreover, our results show that BETA mitigates robustness overfitting and is roughly five times as effective as AutoAttack.
Theoretical underpinnings of surrogate minimization. In this paper, we focused on the empirical performance of AT in the context of the literature concerning adversarial examples in computer vision. However, the study of the efficacy of surrogate losses in minimizing the target 0-1 loss is a well studied topic among theorists. Specifically, this literature considers two notions of minimizers for the surrogate loss also minimizing the target loss: (1) consistency, which requires uniform convergence, and (2) calibration, which requires the weaker notion of pointwise convergence (although (Bartlett et al., 2006) shows that these notions are equivalent for standard, i.e., non-adversarial, classification).
In the particular case of classification in the presence of adversaries, (Bao et al., 2020) and (Meunier et al., 2022) claimed that for the class of linear models, no convex surrogate loss is calibrated with respect to the 0-1 zero-sum formulation of AT, although certain classes of nonconvex losses can maintain calibration for such settings. However, in (Awasthi et al., 2021), the authors challenge this claim, and generalize the calibration results considered by (Bao et al., 2020) beyond linear models. One interesting direction future work would be to provide a theoretical analysis of BETA with respect to the margin-based consistency results proved very recently in (Frank & Niles-Weed, 2023). We also note that in parallel, efforts have been made to design algorithms that are approximately calibrated, leading to—among other things—the TRADES algorithm (Zhang et al., 2019), which we compare to in Section 5. Our work is in the same vein, although BETA does not require approximating a divergence term, which leads to non-calibration of the TRADES objective.
@@ 7 Conclusion
In this paper, we argued that the surrogate-based relaxation commonly employed to improve the tractability of adversarial training voids guarantees on the ultimate robustness of trained classifiers, resulting in weak adversaries and ineffective algorithms. This shortcoming motivated the formulation of a novel, yet natural bilevel approach to adversarial training and evaluation in which the adversary and defender optimize separate objectives, which constitutes a non-zero-sum game. Based on this formulation, we developed a new adversarial attack algorithm (BETA) and a concomitant AT algorithm, which we call BETA-AT. In our experiments, we showed that BETA-AT eliminates robust overfitting and we showed that even when early stopping based model selection is used, BETA-AT performs comparably to AT. Finally, we showed that BETA provides almost identical estimates of robustness to AutoAttack.
@@ Acknowledgements
FL is funded (in part) through a PhD fellowship of the Swiss Data Science Center, a joint venture between EPFL and ETH Zurich. VC is supported by the Hasler Foundation Program: Hasler Responsible AI (project number 21043), the Army Research Office under grant number W911NF-24-1-0048, and the Swiss National Science Foundation (SNSF) under grant number 200021-205011. AR, HH, and GP are supported by the NSF Institute for CORE Emerging Methods in Data Science (EnCORE). AR is also supposed by an ASSET Amazon AWS Trustworthy AI Fellowship.
Appendices
[…]
@@ Appendix A Proof of proposition 1
Suppose that there exists $\hat{\eta}$ satisfying $\left|\left|\hat{\eta}\right|\right|\leq\epsilon$ such that for some $j\in[K]$, $j\neq Y$ we have $M_{\theta}(X+\hat{\eta},Y)_{j}>0$. That is, assume that
@@
and for some $\hat{\eta}$ and some $j$ we have $f_{\theta}(X+\hat{\eta})_{j}>f_{\theta}(X+\hat{\eta})_{Y}$, which implies that $\operatorname*{arg\,max}_{j\in[K]}f_{\theta}(X+\hat{\eta})_{j}\neq Y$. Hence, $\hat{\eta}$ induces a misclassification error, i.e.,
$\hat{\eta}\in\operatorname*{arg\,max}_{\eta:\|\eta\|_{2}\leq\epsilon}\left\{\operatorname*{arg\,max}_{j\in[K]}f_{\theta}(X+\eta)_{j}\neq Y\right\}.$ (25)
In particular, if
then it holds that
$\eta^{\star}\in\operatorname*{arg\,max}_{\eta:\|\eta\|_{2}\leq\epsilon}\left\{\operatorname*{arg\,max}_{j\in[K]}f_{\theta}(X+\eta)_{j}\neq Y\right\}.$ (27)
Otherwise, if it holds that
@@
then for all $\eta:\>\left|\left|\eta\right|\right|<\epsilon$ and all $j\neq Y$, we have $f_{\theta}(X+\eta)_{j}<f_{\theta}(X+\eta)_{Y}$, so that $\operatorname*{arg\,max}_{j\in[K]}f_{\theta}(x+\eta)_{j}=Y$, i.e., there is no adversarial example in the ball. In this case, for any $\eta$, if it holds that
@@ Appendix B Smooth reformulation of the lower level
First, note that the problem in eqs. 21, 22 and 23 is equivalent to
In this formulation, both upper- and lower-level problems are smooth (barring the possible use of nonsmooth components like ReLU). Most importantly (I) the smoothing is obtained through a lower bound of the original objective in eqs. 22 and 23, retaining guarantees that the adversary will increase the misclassification error and (II) all the adversarial perturbations obtained for each class now appear in the upper level (40), weighted by their corresponding negative margin. In this way, we make efficient use of all perturbations generated: if two perturbations from different classes achieve the same negative margin, they will affect the upper-level objective in fair proportion. This formulation gives rise to algorithm 3.
Appendix C Running time analysis
Figure 2: Adversarial training performance-speed trade-off. Each point is annotated with the number of steps with which the corresponding algorithm was run. Observe that robust overfitting is eliminated by BETA, but that this comes at the cost of increased computational overhead. This reveals an expected performance-speed trade-off for our algorithm.
Figure 3: Adversarial evaluation timing comparison. The running time for evaluating the top models on RobustBench using AutoAttack and BETA with the same settings as Table 2 are reported. On average, BETA is 5.11 times faster than AutoAttack.
C.1 Speed-performance trade-off
In Figure 2, we analyze the trade-off between the running time and performance of BETA. Specifically, on the horizontal axis, we plot the running time (in seconds) of an epoch of BETA, and on the vertical axis we plot the performance measured via the robust accuracy with respect to a 20-step PGD adversary. We compare BETA to PGD and TRADES, and we show the speed-performance trade-off when each of these algorithms are run for 5, 10, and 20 iterations; the iteration count is labeled next to each data point. The leftmost panel shows early stopping model selection, and the rightmost panel shows last iterate model selection. Notice that while BETA is significantly more resource intensive than PGD and TRADES, BETA tends to outperform the baselines, particularly if one looks at the gap between early stopping and last iterate model selection.
C.2 Evaluation running time analysis
We next analyze the running time of BETA when used as to adversarially evaluate state-of-the-art robust models. In particular, we return to the setting of Table 2, wherein we compared the performance of AutoAttack to BETA. In Figure 3, we show the wall-clock time of performing adversarial evaluation using both of these algorithms. Notice that AutoAttack takes significantly longer to evaluate each of these models, and as we showed in Table 2, this additional time does not yield a better estimate of the robustness of these models. Indeed, by averaging over the scores in Figure 0(b), we find that BETA is 5.11$\times$ faster than AutoAttack on average.
Appendix D Utility of maximizing the surrogate loss
In this appendix, we show that there exists cases in which our margin-based inner maximization retrieves the optimal adversarial perturbation while the standard inner max with the surrogate loss fails to do so. In this example, we consider a classification problem in which the classifier $f:\mathbb{R}^{2}\to\mathbb{R}^{3}$ is linear across three classes $\{1,2,3\}$. Specifically we define $f$ in the following way:
$f(x_{1},x_{2})=\left[\begin{array}[]{cc}0&-1\\ -1&0\\ 1&0\end{array}\right]\left[\begin{array}[]{c}x_{1}\\ x_{2}\end{array}\right]$ (42)
Furthermore, let $\epsilon=0.8$, let $(x_{1},x_{2})=(0,-1)$, and assume without loss of generality that the correct class is $y=1$. The solution for the maximization of cross-entropy loss is given by:
$\max_{\|\eta\|\leq 0.8}\ell(f(x+\eta),1)=\max_{\|\eta\|\leq 0.8}-\log\left(\dfrac{e^{1-\eta_{2}}}{e^{1-\eta_{2}}+e^{-\eta_{1}}+e^{\eta_{1}}}\right)$ (43)
where $\ell$ denotes the cross entropy loss. Now observe that by the monotonicity of the logarithm function, this problem on the right-hand-side is equivalent to the following problem:
$\min_{\|\eta\|\leq 0.8}\dfrac{e^{1-\eta_{2}}}{e^{1-\eta_{2}}+e^{-\eta_{1}}+e^{\eta_{1}}}$ $=1+\max_{\|\eta\|\leq 0.8}\dfrac{e^{-\eta_{1}}+e^{\eta_{1}}}{e^{1-\eta_{2}}}$ (44)
$=\max_{|\eta_{1}|\leq 0.8}\max_{|\eta_{2}|\leq\sqrt{0.8^{2}-\eta_{1}^{2}}}\dfrac{e^{-\eta_{1}}+e^{+\eta_{1}}}{e^{1-\eta_{2}}}$ (45)
where in the final step we split the problem so that we optimize separately over $\eta_{1}$ and $\eta_{2}$. Observe that the inner problem, for which the numerator is constant, satisfies the following:
$\max_{|\eta_{2}|\leq\sqrt{0.8^{2}-\eta_{1}^{2}}}\dfrac{e^{-\eta_{1}}+e^{+\eta_{1}}}{e^{1-\eta_{2}}}=\min_{|\eta_{2}|\leq\sqrt{0.8^{2}-\eta_{1}^{2}}}e^{1-\eta_{2}}=\min_{|\eta_{2}|\leq\sqrt{0.8^{2}-\eta_{1}^{2}}}1-\eta_{2}$ (46)
As the objective is linear in the rightmost optimization problem, it’s clear that $\eta_{2}^{\star}=\sqrt{0.8^{2}-\eta_{1}^{2}}$. Now returning to (45), we substitute $\eta_{2}^{\star}$ and are therefore left to solve the following problem:
$\max_{|\eta_{1}|\leq 0.8}\dfrac{e^{-\eta_{1}}+e^{\eta_{1}}}{e^{1-\sqrt{0.8^{2}-\eta_{1}^{2}}}}$ $=\max_{|\eta_{1}|\leq 0.8}(e^{-\eta_{1}}+e^{\eta_{1}})e^{\sqrt{0.8^{2}-\eta_{1}^{2}}}$ (47)
$=\max_{0\leq\eta_{1}\leq 0.8}\dfrac{e^{-\eta_{1}}+e^{\eta_{1}}}{e^{1-\sqrt{0.8^{2}-\eta_{1}^{2}}}}$ (48)
where in the final step we used the fact that the objective is symmetric in $\eta_{1}$. By visual inspection, this function achieves its maximum at $\eta^{\star}_{1}=0$ (see Figure 4). Hence, the optimal perturbation obtained via cross-entropy maximization is $\eta^{\star}=(0,0.8)$. Therefore,
$(x_{1},x_{2})+(\eta^{\star}_{1},\eta^{\star}_{2})=(0,-1)+(0,0.8)=(0,-0.2)$
Then, by applying the classifier $f$, we find that
$f(0,-0.2)=\left[\begin{array}[]{cc}0&-1\\ -1&0\\ 1&0\end{array}\right]\left[\begin{array}[]{c}0\\ 0.2\end{array}\right]=\left[\begin{array}[]{c}0.2\\ 0\\ 0\end{array}\right]$ (49)
This shows that the class assigned to this optimally perturbed example is still the correct class $y=1$, i.e., the attacker fails to find an adversarial example.
Figure 4: Plot of function to be maximized in eq. 48. We subtract $y=2.5$ for ease of viewing
In contrast, the main idea in the derivation of the BETA algorithm is to optimize the margins separately for both possible incorrect classes $y=2$ and $y=3$. In particular, for the class $y=2$, BETA solves the following problem:
$\max_{\|\eta\|\leq 0.8}([-1,0]-[0-1])\cdot(x+\eta)$ (50)
The point $\eta=[-0.8,0.8]/\sqrt{2}$ is optimal for this linear problem. On the other hand, for the class $y=3$, BETA solves the following problem:
$\max_{\|\eta\|\leq 0.8}([1,0]-[0-1])\cdot(x+\eta)$ (51)
The point $\eta=[0.8,0.8]/\sqrt{2}$ is optimal for this problem. Observe that both achieve the same value of the margin, so BETA can choose either optimal point; without loss of generality, assume that BETA chooses the second point $\eta^{\star}=[0.8,0.8]/\sqrt{2}$ as the optimal solution. The corresponding classifier takes the following form:
$f(0.8/\sqrt{2},0.8/\sqrt{2}-1)$ $=\left[\begin{array}[]{cc}0&-1\\ -1&0\\ 1&0\end{array}\right]\left[\begin{array}[]{c}0.8/\sqrt{2}\\ 0.8/\sqrt{2}-1\end{array}\right]$ (57)
$=\left[\begin{array}[]{c}1-0.8/\sqrt{2}\\ -0.8/\sqrt{2}\\ 0.8/\sqrt{2}\end{array}\right]$ (61)
$\approx\left[\begin{array}[]{c}0.43\\ -0.57\\ 0.57\end{array}\right]$ (65)
Hence, the classifier returns the incorrect class, i.e., the attack is successful. This shows that whereas the cross-entropy maximization problem fails to find an adversarial example, BETA succeeds in finding an adversarial example.